### v1

# Intertwined Nematicity, Multiferroicity, and Nonlinear Hall Effect in Rhombohedral Pentalayer Graphene

thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.
Intertwined Nematicity, Multiferroicity, and Nonlinear Hall Effect in Rhombohedral Pentalayer Graphene
Erin Morissette1    Peiyu Qin1    K. Watanabe2    T. Taniguchi3    J.I.A. Li1 jia_li@brown.edu 1Department of Physics, Brown University, Providence, RI 02912, USA 2Research Center for Functional Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 3International Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan
(March 13, 2025)
Electronic nematicity-the spontaneous reduction of rotational symmetry-has been widely studied in strongly correlated quantum materials, yet its interplay with other symmetry-breaking phases remains an outstanding experimental challenge to unravel, particularly in van der Waals heterostructures. Here, using angle-resolved transport measurements, we reveal a unique intertwinement of nematicity, orbital multiferroicity, and the nonlinear Hall effect in rhombohedral pentalayer graphene. This coupling enables electric-field-driven switching of the nematic principal axis and the nonlinear Hall polar vector, coinciding with the butterfly-shaped hysteresis of the multiferroic order. We identify two sharply defined nematic phase transitions: the first coincides with the onset of orbital ferromagnetism, while the second, at lower temperatures, aligns with the emergence of ferroelectricity. Our findings point to a nematic-driven instability that intertwines rotational, time-reversal, and inversion symmetry breaking. This is consistent with spontaneous momentum polarization, proposed to occur under the influence of the Coulomb-driven flocking effect. Together, our observations establish nematicity as a fundamental organizing principle in the landscape of intertwined electronic orders in strongly correlated two-dimensional systems.
Figure 1: Rhombohedral-stacking pentalayer graphene and angle-resolved transport measurement. (a) Longitudinal resistance $R_{\parallel}$ as a function of carrier density $n$ and displacement field $D$ in a dual-encapsulated, dual-gated rhombohedral pentalayer graphene device. The inset shows schematic diagram of the heterostructure. (b) Schematic of the angle-resolved transport measurement setup used to extract $R_{\parallel}$ (left) and $R_{\perp}$ (right) in a disk-shaped sample. (c) Angular dependence of $R_{\parallel}$ and $R_{\perp}$ inside the triangular multiferroic region, measured at $T=40$ mK and $B=0$ T. (d) $R_{0}$ and (e) $R_{H}$, extracted from the angular dependence using Eq. 1-2, measured as $B$ is swept back and forth.
Figure 2: Multiferroicity and electric-field-driven switching of transport anisotropy. Parameters describing the angular dependence are extracted from the multiferroic regime. (a) $R_{0}$ and (b-c) $R_{H}$ extracted from angle-resolved transport measurements as a function of $n$ and $D$ around the multiferroic regime. $R_{H}$ exhibits a sign reversal as $D$ is swept (b) from positive to negative, and (c) from negative to positive. As $D$ is swept back and forth at $n=-0.82\times 10^{12}$ cm-2, the multiferroic order manifests in the electric-field-driven switching in (d) the anomalous Hall effect $R_{H}$, which is accompanied by the hysteretic behaviors in (e) $R_{0}$, (f) $\Delta R/R_{0}$, and (g) $\alpha$. (h-i) Angular dependence measured at $n=-0.82\times 10^{12}$ cm-2 and $D=25$ mV/nm as $D$ is swept (h) backwards and (i) forward.
Figure 3: Nematic phase transitions. (a-b) Temperature dependence of (a) the magnitude of the anomalous Hall coefficient $|R_{H}|$ and (b) the strength of transport anisotropy $\Delta R/R_{0}$. Upper right inset shows a zoomed in plot of $\Delta R/R_{0}$ around $T_{nem}^{1}$, where the dashed curve denote the best fit using the Curie-Weiss law. (c-e) Angular dependence of $R_{\parallel}$ and $R_{\perp}$ measured (c) in temperature regime i at $T=10$ K, (d) in temperature regime ii at $T=2.8$ K, and (e) in temperature regime iii at $T=0.4$ K. (f-h) Anomalous Hall coeffecient $R_{H}$ measured with sweeping $D$ back and forth (f)) in temperature regime i at $T=5$ K, (g) in temperature regime ii at $T=3.5$ K, and (h) in temperature regime iii at $T=0.04$ K.
Figure 4: Nonlinear Hall effect. Angular dependence of the second-harmonic nonlinear transport response (a) parallel ($V_{\parallel}^{2\omega}$) and (b) perpendicular ($V_{\perp}^{2\omega}$) to current flow. $\Delta V_{\perp}^{2\omega}$ denotes the angular oscillation in $V_{\perp}^{2\omega}$. (c) Angular dependence of linear transport response measured at the same $n$ and $D$ compared to panels (a) and (b). (d) $\Delta V_{\perp}^{2\omega}$ (top panel) and $|R_{H}|$ (bottom panel) measured as a function of $T$ at $D=10$ mV/nm and $n=-0.58\times 10^{12}$ cm-2. (e) $\Delta V_{\perp}^{2\omega}$ (top panel) and $|R_{H}|$ (bottom panel) measured as a function of $D$ at $T=3.5$ K and $n=-0.58\times 10^{12}$ cm-2. (f-g) The angular dependence of $V_{\perp}^{2\omega}$ measured at the same $n$ and $D$, with different sweeping directions in $D$. (h) Schematic diagram of the polar vector $\mathcal{C}$ extracted from the angular dependence in panels (f) and (g).
Figure 5: Fermiology and coexistence of electron and hole fermi pockets. (a) $R_{\parallel}$ measured as a function of $n$ and $B$ measured at $D=75$ mV/nm. (b) Fourier transform of quantum oscillation in (a). (c) Solid lines label the center frequencies of the most prominent peaks in fν from (b). Inset shows possible fermi surface contours in different density regimes, extracted based on the intensity peaks in fν. (d-e) Fourier transform amplitude of data in (b) for $n=-0.458\times 10^{12}$ cm-2 (d) and $n=-1.325\times 10^{12}$ cm-2 (e).
In strongly interacting quantum materials, the interplay of correlations, broken symmetry, and topology gives rise to a complex landscape of intertwined electronic orders at low temperature  Fradkin et al. (2015); Keimer and Moore (2017); Fernandes et al. (2019); Nuckolls and Yazdani (2024); Wang et al. (2023). Among these, nematicity has been shown to play a pivotal role in shaping the low-temperature phase diagram  Kivelson et al. (1998); Fradkin et al. (2010); Oganesyan et al. (2001); Fernandes et al. (2014); Böhmer et al. (2022); Cao et al. (2020); Zhang et al. (2024). Extensive research has revealed its deep connection to competing phases. In iron-based superconductors, for example, nematicity frequently coexists with magnetic and superconducting transitions, underscoring a fundamental yet still poorly understood connection between these emergent orders  Chu et al. (2012); Fernandes et al. (2014); Böhmer et al. (2022). Despite its ubiquity across material platforms, the precise nature of nematicity remains elusive in most quantum materials, especially low-dimensional systems, primarily due to experimental challenges in isolating and characterizing this interaction-driven order.
This raises several fundamental questions: What constitutes definitive evidence of Coulomb-driven nematicity in transport measurement, distinct from lattice distortions  Fernandes et al. (2014); Böhmer et al. (2022); Ren et al. (2021); Calvera et al. (2024)? Can nematicity be controlled and switched, similar to other Coulomb-driven instabilities, such as ferromagnetism and ferroelectricity  Samajdar et al. (2021); Naritsuka et al. (2023); Klein et al. (2020)? What is the precise relationship between nematicity, other symmetry-breaking orders, and topological phases  Samkharadze et al. (2016); Sun et al. (2009)? Resolving these challenges requires experimental strategies capable of identifying intrinsic nematic order using transport method—a crucial step toward unraveling the mechanisms governing electronic symmetry breaking and enabling new approaches to manipulate emergent quantum phenomena.
In this work, we investigate the low-temperature phase diagram of rhombohedral-stacked pentalayer graphene (rPG), a system where inherently flat energy bands and nontrivial topology lead to a striking interplay between ferromagnetism and ferroelectricity, culminating in the emergence of multiferroicity  Han et al. (2023, 2024); Dong et al. (2024); Lu et al. (2024). Using angle-resolved transport measurement in both the linear and nonlinear regimes  Wu et al. (2017, 2020); Vafek (2023); Chichinadze et al. (2024); Zhang et al. (2024), we establish a previously unrecognized connection between nematicity, the nonlinear Hall effect  Sodemann and Fu (2015); Ma et al. (2019); Kang et al. (2019); Chichinadze et al. (2024), and multiferroicity  Han et al. (2023). This coupling manifests in electric-field-driven switching and hysteretic behaviors in both nematicity and the nonlinear Hall effect, demonstrating unprecedented control over electronic symmetry breaking. Additionally, our results uncover a pair of thermodynamic transitions, driven by nematic instabilities, that intertwine rotational, time-reversal, and inversion symmetry breaking, imposing stringent constraints on the underlying electronic order in rhombohedral graphene.
We begin by investigating the transport properties of rPG in the linear regime, where an ohmic response is probed using a small d.c. current bias of $I_{dc}=5$ nA. Fig. 1a plots the longitudinal resistance $R_{\parallel}$ across the low-temperature phase space, parameterized by charge carrier density ($n$) and displacement field ($D$). The resulting $n-D$ map exhibits high symmetry around $D=0$, consistent with the absence of substrate-induced sublattice symmetry breaking, indicative of misalignment between rPG and hBN. Notably, a triangular region emerges near $D=0$ on the hole-doping side of the charge-neutrality (CNP), hosting an exotic multiferroic order at low temperature  Han et al. (2023). The remainder of this report focuses on unraveling the complex intertwinement of competing orders within this region.
To perform angle-resolved transport measurements  Chichinadze et al. (2024); Zhang et al. (2024), the rPG sample is patterned in a “sunflower” geometry, featuring a disk-shaped area with eight evenly spaced electrical contacts, as illustrated in Fig. 1b. Transport measurements are conducted by applying a current bias between two contacts, designated as source (S) and drain (D), with their relative orientation defining the azimuth direction of current flow ($\phi$). By measuring the voltage response between contacts $A$ and $B$, we extract the longitudinal ($R_{\parallel}$) and transverse ($R_{\perp}$) resistance.
Fig. 1c presents angle-dependent transport responses of the multiferroic state. As a function of varying $\phi$, both longitudinal ($R_{\parallel}$) and transverse ($R_{\perp}$) resistances exhibit a two-fold angular oscillation, described by:
$R_{\parallel}(\phi)$ $=$ $R_{0}-\Delta R\cos 2(\phi-\alpha),$ (1)
$R_{\perp}(\phi)$ $=$ $R_{H}+\Delta R\sin 2(\phi-\alpha).$ (2)
Here, $\Delta R$ represents the oscillation amplitude, $R_{0}$ is the average value of $R_{\parallel}$$(\phi)$, $R_{H}$ is the average value of $R_{\perp}$$(\phi)$, and $\alpha$ denotes the principal axis orientation, defined by the current direction of the highest conductivity. According to theoretical discussion  Vafek (2023), Eq. 1-2 fully determines the conductivity matrix describing transport response in the linear regime: the ratio between maximum and minimum conductivity, qualified by $\Delta R/R_{0}$, offers a direct measure of anisotropy strength. Furthermore, a non-zero $R_{H}$ indicates the presence of antisymmetric off-diagonal terms in the conductivity matrix, corresponding to an anomalous Hall coefficient associated with orbital ferromagnetism.
[…]
In addition to its coupling to magnetism, the anomalous Hall effect exhibits electric-field-driven switching, a hallmark of multiferroic order Han et al. (2023). Figures 2a–c show that a large anomalous Hall coefficient ($R_{H}$) emerges in the low-density portion of the triangular region, with red and blue regions indicating opposite Hall responses. A sign reversal in $R_{H}$ occurs across $D=0$, suggesting that the Hall response is controlled by the displacement field. Notably, $R_{H}$ is highly sensitive to the direction of $D$ sweeping: at a fixed point in the $n$$D$ phase space, reversing the $D$ sweep direction induces a corresponding sign reversal in $R_{H}$, giving rise to an electric-field-driven hysteresis in the anomalous Hall effect.
Most remarkably, this electric-field-driven switching is directly linked to rotational symmetry breaking. To demonstrate this connection, we perform angle-resolved transport along the vertical dashed lines in Fig. 2a–c, extracting the anisotropy parameters from Eqs. 1–2. As $D$ is swept back and forth, $R_{H}$ exhibits a butterfly-shaped hysteresis loop (Fig. 2d), characteristic of the multiferroic nature. Interestingly, both $\Delta R/R_{0}$ and $\alpha$ exhibit hysteresis behaviors that coincide with the electric-field-driven switching.
To further elucidate this switching, Fig. 2h–i compare the angular dependence of transport measured at $D=25$ mV/nm, correspondng to the red and blue open circles in Fig. 2d-g. When $D$ is swept backward (from positive to negative), the transport response exhibits a large angular oscillation, indicating strong anisotropy, accompanied by a positive $R_{H}$. In contrast, as $D$ is swept forward (from negative to positive), a reduced $\Delta R/R_{0}$ signals a suppression of transport anisotropy, while the principal axis ($\alpha$) undergoes a significant rotation. Notably, at negative $D$, strong (weak) anisotropy occurs with forward (backward) $D$-sweeping (see Fig. XX), opposite to the behavior observed in Fig. 2h–i, indicating a strong coupling between electronic anisotropy and the ferroelectric switching.
This hysteretic switching in transport anisotropy, driven by an electric-field, demonstrates unprecedented control over the emergence of rotational symmetry breaking. Such control is fundamentally distinct from the effects of strain and distortion in the crystal lattice, suggesting that the observed transport anisotropy is primarily of electronic origin.
Furthermore, the electric-field-driven switching also underscores the intrinsic connection between broken rotational symmetry, valley polarization, and layer polarization. Due to the abrupt changes in both $\Delta R/R_{0}$ and $\alpha$, the transport response measured along certain azimuth angles does not fully capture the electric-field-driven transitions. For instance, along $\phi=22.5^{\circ}$, $R_{\perp}$ does not exhibit the characteristic butterfly-shaped hysteresis (see Fig. M3). In this context, angle-resolved transport serves as an essential tool for unraveling the unique intertwinement between nematicity and multiferroicity.
To further disentangle these intertwined orders, we analyze their temperature dependence. Fig. 3a-b plot the temperature dependence of $|R_{H}|$ and $\Delta R/R_{0}$, extracted from angle-resolved transport measurements (see Method Section for measurement details). This analysis reveals three distinct temperature regimes, separated by two transitions, denoted as $T_{\text{nem}^{1}}$ and $T_{\text{nem}^{2}}$, respectively.
  • Regime (i)-Isotropic transport: At elevated temperature, angle-resolved transport exhibits a largely isotropic response with a vanishing Hall effect (Fig. 3c).
  • Regime (ii)-Onset of anomalous Hall and anisotropy: As temperature decreases, an anomalous Hall effect emerges, with a sharp onset in $|R_{H}|$ around $T=4$ K (Fig. 3a), accompanied by an enhanced transport anisotropy (Fig. 3b and d). Notably, ferroelectricity is absent in this regime, as the anomalous Hall effect does not exhibit electric-field-driven switching under sweeping $D$ (Fig. 3g).
  • Regime (iii)-Emergence of Multiferroicity: Below $T=1.5$ K, the butterfly-shaped hysteresis in $R_{H}$ (Fig. 3h) signals the emergence of multiferroicity, which coincides with a second sharp increase in transport anisotropy (Fig. 3b).
The temperature dependence of $\Delta R/R_{0}$ just above $T_{\text{nem}^{1}}$ is well approximated by a Curie-Weiss curve, diverging at $T_{0}=2.1$ K (inset in Fig. 3b). This behavior closely resembles nematic phase transitions observed in other quantum materials, such as Iron-based superconductors  Chu et al. (2012); Fernandes et al. (2014); Böhmer et al. (2022). This temperature dependence provides strong evidence for a thermodynamic transition driven by a nematic instability, marking a previously unobserved phenomenon in 2D van der Waals materials.
Furthermore, the simultaneous emergence of nematicity and orbital ferromagnetism at $T_{\text{nem}}^{1}$ suggests a common underlying mechanism that intertwines rotational and time-reversal symmetry breaking. This behavior aligns with theoretical predictions of Coulomb-driven instabilities in momentum space, where a collective flocking effect induces spontaneous momentum condensation  Dong et al. (2023); Jung et al. (2015); Huang et al. (2023); Mandal and Fernandes (2023); Parra-Martinez et al. (2025).
Previous theoretical studies have suggested that momentum polarization can manifest through the nonlinear Hall effect  Huang et al. (2023). Here, we provide experimental evidence demonstrating a direct connection between the anomalous Hall effect, nematicity, and the nonlinear Hall effect.
The nonlinear transport response is extracted by measuring the voltage response at the second harmonic frequency of an AC current bias  Ma et al. (2019); Kang et al. (2019); He et al. (2022); Zhang et al. (2024); Chichinadze et al. (2024). To ensure measurements remain within the quadratic response regime, we apply a small AC current of $10-15$ nA (see Methods section for details). Fig. 4a-b present the angular dependence of the nonlinear response inside the multiferroic regime, revealing hallmark signatures of the nonlinear Hall effect. Both the parallel ($V_{\parallel}^{2\omega}$) and perpendicular ($V_{\perp}^{2\omega}$) channels of the nonlinear response exhibit a one-fold angular oscillation, with $V_{\perp}^{2\omega}$ displaying a significantly larger oscillation amplitude. This behavior is in excellent agreement with the predicted angular response of the nonlinear Hall effect  Chichinadze et al. (2024). Most remarkably, our findings indicate that the nonlinear Hall component plays a dominating role in the conductivity tensor describing transport response in the nonlinear regime.
The nonlinear Hall effect is characterized by a polar vector $\mathcal{C}$, defined by the direction of maximum $V_{\perp}^{2\omega}$ (indicated by green solid lines in Fig. 4a–b). Notably, $\mathcal{C}$ is parallel to the principal axis of maximum resistivity (minimum conductivity) in the linear transport regime (Fig. 4c), supporting a common origin for nematicity and the nonlinear Hall effect.
The nonlinear Hall effect also exhibits a strong correlation with orbital ferromagnetism, as evidenced by the direct link between its magnitude—characterized by the oscillation amplitude of $V_{\perp}^{2\omega}$ ($\Delta V_{\perp}^{2\omega}$)—and the anomalous Hall coefficient ($R_{H}$). The emergence of the nonlinear Hall effect, identified by the onset of $\Delta V_{\perp}^{2\omega}$ as temperature decreases, aligns precisely with the onset of the anomalous Hall effect (Fig. 4d). Furthermore, as shown in Fig. 4d, at a fixed temperature and carrier density, $\Delta V_{\perp}^{2\omega}$ reaches its maximum away from $D=0$, forming two distinct peaks that coincide with maxima in $R_{H}$.
Taken together, the observed correlation between nematicity, the nonlinear Hall effect, and orbital ferromagnetism provides unambiguous evidence for a Coulomb-driven instability that intertwines rotational, time-reversal, and inversion symmetry breaking. This phenomenon is in excellent agreement with the theoretical framework of spontaneous momentum polarization.
Having examined the nature of $T_{\text{nem}}^{1}$, we now turn to the low-temperature transition at $T_{\text{nem}}^{2}$, which coincides with the emergence of ferroelectricity. The fact that the anomalous Hall effect and electric-field-driven switching emerge at distinct temperatures suggests that ferromagnetism and ferroelectricity originate from different mechanisms, despite occupying the same region of the $n$$D$ phase space at low temperatures.
Furthermore, the coincidence of electric-field-driven hysteresis with the sharp onset in $\Delta R/R_{0}$ provides compelling evidence that ferroelectric order is intrinsically anisotropic, a crucial property not accounted for in previous theoretical models of multiferroicity in this system Das and Huang (2024).
The interplay between ferroelectricity and rotational symmetry breaking is further reflected in the tunability of the nonlinear Hall effect at $T<T_{nem}^{2}$. Figures 4g–h plot the angular dependence of $V_{\perp}^{2\omega}$, measured at a fixed carrier density ($n$) and displacement field ($D$), revealing a prominent rotation of the nonlinear Hall vector ($\mathcal{C}$) as $D$ is swept in opposite directions (Fig. 4f). This $D$-induced hysteresis not only establishes an unprecedented degree of control over the nonlinear Hall effect, but also underscores an intrinsic link with nematicity and multiferroicity.
Lastly, we investigate the fermiology of the multiferroic state by analyzing the Shubnikov–de Haas (SdH) quantum oscillations. As shown in Fig. 5a, quantum oscillations exhibit a “fish-net” pattern across a wide carrier density range. We perform a Fourier transform of $R_{\parallel}$ as a function of $1/B$ to identify the the Luttinger volume and degeneracy of the Fermi sea responsible for the “fish-net” pattern (Fig. 5b).
Figure 5c marks the center frequencies of the most prominent peaks in the Fourier spectrum, revealing distinct density regimes characterized by different Fermi surface topologies. At low carrier density ($n=-0.5\times 10^{12}$ cm-2), the two dominant frequencies, $f_{1}$ and $f_{2}$, satisfy $f_{1}-f_{2}=1$, suggesting two possible Fermi surface topologies: (i) a quarter metal phase characterized by an annular Fermi sea or (ii) an exotic half-metallic state, where an electron pocket coexists with a hole pocket, each featuring opposite valley polarizations (inset of Fig. 5c).
[…]
Notably, angle-resolved transport measurements reveal no anomalous Hall effect in the low- and high-density regimes (Fig. 2b-c). This is inconsistent with a quarter metal phase at low-density, suggesting that the Fermi surface instead consists of coexisting electron and hole pockets with opposite valley polarizations, where the overall Chern number remains close to zero. The universal presence of electron and hole pockets is in line with prior theoretical predictions Das and Huang (2024). In the presence of an out-of-plane electric field, the electron and hole pockets polarize to opposite graphene layers, creating a distinct charge distribution that likely serves as the direct link between nematicity and ferroelectricity.
During the transition between these high- and low-density regimes, the spectral weight of the Fourier transform shifts from $f_{1}$ and $f_{2}$ to $f_{3}$ and $f_{4}$, signaling the onset of a partially isospin-polarized (PIP) phase  Zhou et al. (2021); Holleis et al. (2025). Notably, the emergence of multiferroicity at $B=0$ coincides with the density range of the PIP phase, suggesting that Fermi surface reconstruction plays a pivotal role in driving momentum-space instabilities. These instabilities, in turn, are responsible for the simultaneous breaking of time-reversal, rotational, and inversion symmetries.
According to prior theoretical discussions, spontaneous momentum polarization could arise from the instability of an annular Fermi sea  Huang et al. (2023); Jung et al. (2015), or from exchange-driven condensation between discrete Fermi pockets  Dong et al. (2023); Parra-Martinez et al. (2025). These mechanisms provide a compelling framework for understanding the interplay between Fermi surface topology, isospin physics, and emergent ferroic orders in rhombohedral graphene.
Overall, through angle-resolved transport measurements, we uncover the simultaneous emergence of nematicity, multiferroicity, and the nonlinear Hall effect, establishing an unprecedented scenario where rotational, time-reversal, and inversion symmetries are broken in tandem. These findings provide strong experimental support for the Coulomb-driven phenomenon of spontaneous momentum-space condensation  Huang et al. (2023); Jung et al. (2015); Dong et al. (2023); Parra-Martinez et al. (2025), which has been theoretically discussed in connection with the exotic chiral superconductivity.
acknowledgments
J.I.A.L. wishes to acknowledge helpful discussion with Chunli Huang and Rafael Fernandes. E.M. and J.I.A.L. acknowledge support from U.S. National Science Foundation under Award DMR-2143384. J.I.A.L. acknowledge partial support from the Air Force Office of Scientific Research. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. Part of this work was enabled by the use of pyscan (github.com/sandialabs/pyscan), scientific measurement software made available by the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy.