### v2

# Coulomb-driven Momentum Space Condensation in Rhombohedral Hexalayer Graphene

thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.
Coulomb-driven Momentum Space Condensation in Rhombohedral Hexalayer Graphene
Erin Morissette1    Peiyu Qin1    K. Watanabe2    T. Taniguchi3    J.I.A. Li1 jia_li@brown.edu 1Department of Physics, Brown University, Providence, RI 02912, USA 2Research Center for Functional Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 3International Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan
(August 8, 2025)
In systems of strongly correlated electrons, the emergence of spontaneous symmetry breaking serves as a defining hallmark of the underlying many-body physics. Here, we report the discovery of an unprecedented emerging order in rhombohedral hexalayer graphene, defined by the intertwinement of rotational, time-reversal, and inversion symmetry breaking. The link between these broken symmetries is uncovered through angle-resolved transport measurements. The spontaneous breaking of rotational symmetry is identified by the Curie-Weiss behavior in the temperature-dependent onset of transport anisotropy. Most remarkably, the nematic transition, defined by the cusp of transport anisotropy, coincides with the emergence of the anomalous Hall effect in both the linear and nonlinear regimes. Our findings provide unambiguous experimental evidence of spontaneous condensation in momentum space, a phenomenon previously proposed to arise from a Coulomb-driven flocking effect. Together, these observations establish electronic anisotropy as a fundamental organizing principle in the landscape of intertwined electronic orders in strongly correlated two-dimensional systems.
Figure 1: Rhombohedral-stacking hexalayer graphene and angle-resolved transport measurement. (a) Longitudinal resistance $R_{\parallel}$ as a function of carrier density $n$ and displacement field $D$ in a dual-encapsulated, dual-gated rhombohedral hexalayer graphene device. The inset shows schematic diagram of the heterostructure. (b) Schematic of the angle-resolved transport measurement setup used to extract $R_{\parallel}$ (left) and $R_{\perp}$ (right) in a disk-shaped sample. (c) Angular dependence of $R_{\parallel}$ and $R_{\perp}$ inside the triangular multiferroic region, measured at $T=40$ mK and $B=0$ T. (d) $R_{0}$ and (e) $R_{H}$, extracted from the angular dependence using Eq. 1-2, measured as $B$ is swept back and forth.
Figure 2: Multiferroicity and electric-field-driven switching of transport anisotropy. Parameters describing the angular dependence are extracted from the multiferroic regime. (a) $R_{0}$ and (b-c) $R_{H}$ extracted from angle-resolved transport measurements as a function of $n$ and $D$ around the multiferroic regime. $R_{H}$ exhibits a sign reversal as $D$ is swept (b) from positive to negative, and (c) from negative to positive. As $D$ is swept back and forth at $n=-0.82\times 10^{12}$ cm-2, the multiferroic order manifests in the electric-field-driven switching in (d) the anomalous Hall effect $R_{H}$, which is accompanied by the hysteretic behaviors in (e) $R_{0}$, (f) $\Delta R/R_{0}$, and (g) $\alpha$. (h-i) Angular dependence measured at $n=-0.82\times 10^{12}$ cm-2 and $D=25$ mV/nm as $D$ is swept (h) backwards and (i) forward.
Figure 3: Nematic phase transitions. (a-b) Temperature dependence of (a) the magnitude of the anomalous Hall coefficient $|R_{H}|$ and (b) the strength of transport anisotropy $\Delta R/R_{0}$. Upper right inset shows a zoomed in plot of $\Delta R/R_{0}$ around $T_{nem}^{1}$, where the dashed curve denote the best fit using the Curie-Weiss law. (c-e) Angular dependence of $R_{\parallel}$ and $R_{\perp}$ measured (c) in temperature regime i at $T=10$ K, (d) in temperature regime $i$ at $T=2.8$ K, and (e) in temperature regime iii at $T=0.4$ K. (f-h) Anomalous Hall coeffecient $R_{H}$ measured with sweeping $D$ back and forth (f)) in temperature regime i at $T=5$ K, (g) in temperature regime $i$ at $T=3.5$ K, and (h) in temperature regime iii at $T=0.04$ K.
Figure 4: Nonlinear Hall effect. Angular dependence of the second-harmonic nonlinear transport response (a) parallel ($V_{\parallel}^{2\omega}$) and (b) perpendicular ($V_{\perp}^{2\omega}$) to current flow. $\Delta V_{\perp}^{2\omega}$ denotes the angular oscillation in $V_{\perp}^{2\omega}$. (c) Angular dependence of linear transport response measured at the same $n$ and $D$ compared to panels (a) and (b). (d) $\Delta V_{\perp}^{2\omega}$ (top panel) and $|R_{H}|$ (bottom panel) measured as a function of $T$ at $D=10$ mV/nm and $n=-0.58\times 10^{12}$ cm-2. (e) $\Delta V_{\perp}^{2\omega}$ (top panel) and $|R_{H}|$ (bottom panel) measured as a function of $D$ at $T=3.5$ K and $n=-0.58\times 10^{12}$ cm-2. (f-g) The angular dependence of $V_{\perp}^{2\omega}$ measured at the same $n$ and $D$, with different sweeping directions in $D$. (h) Schematic diagram of the polar vector $\mathcal{C}$ extracted from the angular dependence in panels (f) and (g).
Figure 5: Fermiology and coexistence of electron and hole fermi pockets. (a) $R_{\parallel}$ measured as a function of $n$ and $B$ measured at $D=75$ mV/nm. (b) Fourier transform of quantum oscillation in (a). (c) Solid lines label the center frequencies of the most prominent peaks in fν from (b). Inset shows possible fermi surface contours in different density regimes, extracted based on the intensity peaks in fν. (d-e) Fourier transform amplitude of data in (b) for $n=-0.458\times 10^{12}$ cm-2 (d) and $n=-1.325\times 10^{12}$ cm-2 (e).
Rhombohedral multilayer graphene systems have recently emerged as a compelling platform for exploring strongly correlated quantum phenomena  Zhou et al. (2021); Han et al. (2024a); Chen et al. (2020); Han et al. (2023); Lu et al. (2024); Choi et al. (2025); Liu et al. (2024); Sha et al. (2024); Han et al. (2024b). While the energy bands remain dispersive at high carrier densities, the application of a perpendicular electric field ($D$) dramatically enhances the density of states near the band edges, producing a broad range of interaction-driven phases. The resulting low-temperature phase space hosts a diverse spectrum of emergent orders, many of which exhibit spontaneous time-reversal symmetry breaking—including half- and quarter-metallic states  Zhou et al. (2021); Han et al. (2024a); Chen et al. (2020), multiferroicity  Han et al. (2023), and various quantum Hall effect states at zero external magnetic field  Lu et al. (2024); Choi et al. (2025); Liu et al. (2024); Sha et al. (2024); Han et al. (2024b).
Beyond time-reversal symmetry breaking, another critical component of the underlying electronic order has been proposed in several theoretical works  Jung et al. (2015); Huang et al. (2023); Dong et al. (2023). These studies suggest that the quarter-metal phase is accompanied by spontaneous breaking of $C_{3}$ rotational symmetry. Driven by a Coulomb-induced flocking instability, charge carriers spontaneously condense into a single corner of the trigonally-warped Fermi surface, acquiring a finite net momentum. This mechanism links broken time-reversal and rotational symmetries. The resulting electronic ground state, described by momentum polarization, may play a foundational role in shaping the broader phase diagram of rhombohedral multilayer graphene.
Despite its theoretical appeal, direct experimental identification of a momentum-polarized ground state has remained elusive, largely due to the intrinsic difficulty of detecting spontaneous rotational symmetry breaking using conventional probes  Zhang et al. (2024); Lin et al. (2023). Efforts to characterize electronic orders with such broken symmetries raise several fundamental questions: What constitutes definitive evidence of spontaneous rotational symmetry breaking, distinct from lattice-induced distortions  Fernandes et al. (2014); Böhmer et al. (2022); Ren et al. (2021); Calvera et al. (2024)? Can this broken symmetry be externally controlled and switched, in analogy to other Coulomb-driven instabilities such as ferromagnetism Samajdar et al. (2021); Naritsuka et al. (2023); Klein et al. (2020)? And finally, what are the distinct transport signatures of an electronic state that simultaneously breaks time-reversal and rotational symmetries  Huang et al. (2023)?
Motivated by these questions, in this work we investigate the low-temperature phase diagram of rhombohedral-stacked hexalayer graphene (RHG). Using angle-resolved transport measurements in both the linear and nonlinear regimes Wu et al. (2017, 2020); Vafek (2023); Chichinadze et al. (2024); Zhang et al. (2024), we uncover a unique intertwinement between transport anisotropy, the nonlinear Hall effect Sodemann and Fu (2015); Ma et al. (2019); Kang et al. (2019); Chichinadze et al. (2024), and orbital multiferroicity Han et al. (2023). Together, these intertwined orders provide unambiguous experimental support for a momentum-polarized state that simultaneously breaks time-reversal, rotational, and inversion symmetries, offering new insight into the nature of Coulomb-driven symmetry breaking in rhombohedral multilayer graphene.
We begin by investigating the transport properties of RHG in the linear regime, where an ohmic response is probed using a small d.c. current bias of $I_{dc}=5$ nA. Fig. 1a plots the longitudinal resistance $R_{\parallel}$ across the low-temperature phase space, parameterized by charge carrier density ($n$) and displacement field ($D$). The resulting nDn-D map exhibits high symmetry around $D=0$, consistent with the absence of substrate-induced sublattice symmetry breaking, indicative of misalignment between RHG and hBN. Notably, a triangular region emerges near $D=0$ on the hole-doping side of the charge-neutrality (CNP), hosting an exotic multiferroic order at low temperature  Han et al. (2023). The remainder of this report focuses on unraveling the nature of spontaneous symmetry breaking within this region using angle-resolved transport.
To perform angle-resolved transport measurements  Chichinadze et al. (2024); Zhang et al. (2024), the RHG sample is patterned in a “sunflower” geometry, featuring a disk-shaped area with eight evenly spaced electrical contacts, as illustrated in Fig. 1b. Transport measurements are conducted by applying a current bias between two contacts, designated as source (S) and drain (D), with their relative orientation defining the azimuth direction of current flow ($\phi$). By measuring the voltage response between contacts $A$ and $B$, we extract the longitudinal ($R_{\parallel}$) and transverse ($R_{\perp}$) resistance.
Fig. 1c presents angle-dependent transport responses of the multiferroic state. As a function of varying $\phi$, both longitudinal ($R_{\parallel}$) and transverse ($R_{\perp}$) resistances exhibit a two-fold angular oscillation, described by:
$R_{\parallel}(\phi)$ $=$ $R_{0}-\Delta R\cos 2(\phi-\alpha),$ (1)
$R_{\perp}(\phi)$ $=$ $R_{H}+\Delta R\sin 2(\phi-\alpha).$ (2)
Here, $\Delta R$ represents the oscillation amplitude, $R_{0}$ is the average value of $R_{\parallel}$$(\phi)$, $R_{H}$ is the average value of $R_{\perp}$$(\phi)$, and $\alpha$ denotes the principal axis orientation, defined by the current direction of the highest conductivity. According to theoretical discussion  Vafek (2023), Eq. 1-2 fully determines the conductivity matrix describing transport response in the linear regime: the ratio between the oscillation amplitude and average value, $\Delta R/R_{0}$, offers a direct measure of anisotropy strength. Furthermore, a non-zero $R_{H}$ indicates the presence of antisymmetric off-diagonal terms in the conductivity matrix, corresponding to an anomalous Hall coefficient associated with orbital ferromagnetism.
[…]
In addition to its coupling to magnetism, the anomalous Hall effect exhibits electric-field-driven switching, a hallmark of multiferroic order Han et al. (2023). Figures 2a–c show that a large anomalous Hall coefficient ($R_{H}$) emerges in the low-density portion of the triangular region, with red and blue indicating opposite Hall responses. A sign reversal in $R_{H}$ occurs across $D=0$, suggesting that the Hall response is controlled by the displacement field. Notably, $R_{H}$ is highly sensitive to the direction of $D$ sweeping: at a fixed point in the nnDD phase space, reversing the $D$ sweep direction induces a corresponding sign reversal in $R_{H}$, giving rise to an electric-field-driven hysteresis in the anomalous Hall effect.
Most remarkably, this electric-field-driven switching is directly linked to rotational symmetry breaking. To demonstrate this connection, we perform angle-resolved transport along the vertical dashed lines in Fig. 2a–c, extracting the anisotropy parameters from Eqs. 1–2. As $D$ is swept back and forth, $R_{H}$ exhibits a butterfly-shaped hysteresis loop (Fig. 2d), characteristic of the multiferroic nature. Interestingly, both $\Delta R/R_{0}$ and $\alpha$ exhibit hysteresis behaviors that coincide with the electric-field-driven switching.
To further elucidate this switching, Fig. 2h–i compare the angular dependence of transport measured at $D=25$ mV/nm, corresponding to the red and blue open circles in Fig. 2d-g. When $D$ is swept backward (from positive to negative), the transport response exhibits a large angular oscillation, indicating strong anisotropy, accompanied by a positive $R_{H}$. In contrast, as $D$ is swept forward (from negative to positive), a reduced $\Delta R/R_{0}$ signals a suppression of transport anisotropy, while the principal axis ($\alpha$) undergoes a significant rotation. Notably, at negative $D$, strong (weak) anisotropy occurs with forward (backward) $D$-sweeping (see Fig. M1), opposite to the behavior observed in Fig. 2h–i, indicating a strong coupling between electronic anisotropy and the ferroelectric switching.
The hysteretic switching in transport anisotropy, driven by an electric-field, demonstrates an unprecedented level of control over rotational symmetry breaking. Unlike strain- or distortion-induced effects in the crystal lattice, this control arises purely from electronic interactions, indicating that the observed anisotropy is fundamentally of electronic origin.
Because both $\Delta R/R_{0}$ and $\alpha$ undergo abrupt changes across the transition, transport responses measured along a fixed azimuth angle may not fully capture the electric-field-driven symmetry breaking. For example, at $\phi=22.5^{\circ}$, $R_{\perp}$ does not display the characteristic butterfly-shaped hysteresis (see Fig. M3). In this context, angle-resolved transport measurements are essential for uncovering the symmetry and nature of the underlying electronic order in phases with strong transport anisotropy.
Next, we analyze the temperature-dependent onset of rotational and time-reversal symmetry breaking. Fig. 3a-b plot the temperature dependence of $|R_{H}|$ and $\Delta R/R_{0}$, extracted from angle-resolved transport measurements. This analysis reveals three distinct temperature regimes, separated by two transitions, denoted as $T_{\text{nem}^{1}}$ and $T_{\text{nem}^{2}}$, respectively.
  • Regime (i)-Isotropic transport: At elevated temperature, angle-resolved transport exhibits a largely isotropic response with a vanishing Hall effect (Fig. 3c).
  • Regime (ii)-Onset of anomalous Hall and anisotropy: As temperature decreases, an anomalous Hall effect emerges, with a sharp onset in $|R_{H}|$ around $T=4$ K (Fig. 3a), accompanied by an enhanced transport anisotropy (Fig. 3b and d). Notably, ferroelectricity is absent in this regime, as the anomalous Hall effect does not exhibit electric-field-driven switching under sweeping $D$ (Fig. 3g).
  • Regime (iii)-Emergence of Multiferroicity: Below $T=1.5$ K, the butterfly-shaped hysteresis in $R_{H}$ (Fig. 3h) signals the emergence of multiferroicity, which coincides with a second sharp increase in transport anisotropy (Fig. 3b).
The temperature dependence of $\Delta R/R_{0}$ just above $T_{\text{nem}^{1}}$ is well approximated by a Curie-Weiss curve (black dashed line in the inset of Fig. 3b), diverging at $T_{0}=2.1$ K. This behavior closely resembles nematic phase transitions observed in other quantum materials, such as Iron-based superconductors  Chu et al. (2012); Fernandes et al. (2014); Böhmer et al. (2022). This temperature dependence provides strong support for the presence of a nematic instability, marking a previously unobserved phenomenon in 2D van der Waals materials.
Furthermore, the emergence of the anomalous Hall effect at $T_{\text{nem}}^{1}$ suggests a common underlying mechanism that intertwines rotational and time-reversal symmetry breaking. This behavior aligns with theoretical predictions of Coulomb-driven instabilities in momentum space, where a collective flocking effect induces spontaneous momentum condensation  Dong et al. (2023); Jung et al. (2015); Huang et al. (2023); Mandal and Fernandes (2023); Parra-Martinez et al. (2025).
Previous theoretical studies have suggested that momentum polarization can manifest through the nonlinear Hall effect  Huang et al. (2023). Here, we provide experimental evidence demonstrating a direct connection between the anomalous Hall effect, nematicity, and the nonlinear Hall effect.
The nonlinear transport response is extracted by measuring the voltage response at the second harmonic frequency of an AC current bias  Ma et al. (2019); Kang et al. (2019); He et al. (2022); Zhang et al. (2024); Chichinadze et al. (2024). To ensure measurements remain within the quadratic response regime, we apply a small AC current of 101510-15 nA (see Methods section for details). Fig. 4a-b present the angular dependence of the nonlinear response inside the multiferroic regime, revealing hallmark signatures of the nonlinear Hall effect. Both the parallel ($V_{\parallel}^{2\omega}$) and perpendicular ($V_{\perp}^{2\omega}$) channels of the nonlinear response exhibit a one-fold angular oscillation, with $V_{\perp}^{2\omega}$ displaying a significantly larger oscillation amplitude. This behavior is in excellent agreement with the predicted angular response of the nonlinear Hall effect  Chichinadze et al. (2024). Most remarkably, our findings indicate that the nonlinear Hall component plays a dominating role in the conductivity tensor describing transport response in the nonlinear regime.
The nonlinear Hall effect is characterized by a polar vector $\mathcal{C}$, defined by the direction of maximum $V_{\perp}^{2\omega}$ (indicated by green solid lines in Fig. 4a–b). Notably, $\mathcal{C}$ is parallel to the principal axis of maximum resistivity (minimum conductivity) in the linear transport regime (Fig. 4c), supporting a common origin for nematicity and the nonlinear Hall effect.
The nonlinear Hall effect also exhibits a strong correlation with orbital ferromagnetism, as evidenced by the direct link between its magnitude—characterized by the oscillation amplitude of $V_{\perp}^{2\omega}$ ($\Delta V_{\perp}^{2\omega}$)—and the anomalous Hall coefficient ($R_{H}$). The emergence of the nonlinear Hall effect, identified by the onset of $\Delta V_{\perp}^{2\omega}$ as temperature decreases, aligns precisely with the onset of the anomalous Hall effect (Fig. 4d). Furthermore, as shown in Fig. 4d, at a fixed temperature and carrier density, $\Delta V_{\perp}^{2\omega}$ reaches its maximum away from $D=0$, forming two distinct peaks that coincide with maxima in $R_{H}$.
Taken together, the observed correlation between transport anisotropy, the nonlinear Hall effect, and orbital ferromagnetism provides unambiguous evidence for a Coulomb-driven instability that intertwines rotational, time-reversal, and inversion symmetry breaking. This phenomenon is in excellent agreement with the theoretical framework of spontaneous momentum polarization.
Further decreasing temperature unlocks another transition at $T_{\text{nem}}^{2}$, which coincides with the emergence of ferroelectricity. The fact that the anomalous Hall effect and electric-field–driven switching emerge at distinct temperatures suggests that ferromagnetism and ferroelectricity originate from different mechanisms, despite occupying the same region of the nnDD phase space at low temperatures.
At the same time, the coincidence of electric-field–driven hysteresis with the sharp onset in $\Delta R/R_{0}$ provides compelling evidence that ferroelectric order is intrinsically anisotropic, a crucial property not accounted for in previous theoretical models of multiferroicity in this system Das and Huang (2024).
The presence of ferroelectricity at $T<T_{\text{nem}}^{2}$ also gives rise to electric-field–driven tunability in the nonlinear Hall effect. Figures 4g–h plot the angular dependence of $V_{\perp}^{2\omega}$, measured at a fixed carrier density ($n$) and displacement field ($D$). As $D$ is swept in opposite directions, angle-resolved measurement of the nonlinear response uncovers a rotation of $90^{\circ}$ in the nonlinear Hall vector ($\mathcal{C}$) (Fig. 4f). Such a prominent rotation suggests that momentum polarization is decoupled from sample details, such as lattice-induced distortion, allowing it to be freely controlled using the interplay with ferroelectricity.
Lastly, we investigate the fermiology of the multiferroic state by analyzing the Shubnikov–de Haas (SdH) quantum oscillations. As shown in Fig. 5a, quantum oscillations exhibit a “fish-net” pattern across a wide carrier density range. We perform a Fourier transform of $R_{\parallel}$ as a function of $1/B$ to identify the the Luttinger volume and degeneracy of the underlying Fermi sea (Fig. 5b).
Figure 5c marks the center frequencies of the most prominent peaks in the Fourier spectrum, revealing distinct density regimes characterized by different Fermi surface topologies. At low carrier density ($n=-0.5\times 10^{12}$ cm-2), the two dominant frequencies, $f_{1}$ and $f_{2}$, satisfy $f_{1}-f_{2}=1$, suggesting two possible Fermi surface topologies: (i) a quarter metal characterized by an annular Fermi sea or (ii) an exotic half-metal, where an electron pocket coexists with a hole pocket, each featuring opposite valley polarizations (inset of Fig. 5c).
At higher densities ($2.5\times 10^{12}$ $^{-2}<n<-1.0\times 10^{12} cm-$2), the sum rule $f_{3}-f_{4}=1/2$ suggests either (i) a half-metal phase with an annular Fermi sea or (ii) the coexistence of two electron and two hole pockets, indicative of a more complex isospin texture.
Since angle-resolved transport measurements reveal no anomalous Hall effect in the low- and high-density regimes (Fig. 2b-c), the Fermi surface likely consists of coexisting electron and hole pockets with opposite valley polarizations, where the overall Chern number remains close to zero. This is in line with prior theoretical predictions Das and Huang (2024). In the presence of an out-of-plane electric field, the electron and hole pockets polarize to opposite graphene layers, creating a distinct charge distribution that likely serves as the direct link between transport anisotropy and ferroelectricity.
As the multiferroic order emerges during the transition between these high- and low-density regimes, the spectral weight of the Fourier transform shifts from $f_{1}$ and $f_{2}$ to $f_{3}$ and $f_{4}$, signaling the onset of a partially isospin-polarized (PIP) phase Zhou et al. (2021); Holleis et al. (2025). This suggests that Fermi surface reconstruction plays a pivotal role in enabling momentum-space instabilities, which, in turn, are responsible for the simultaneous breaking of time-reversal, rotational, and inversion symmetries.
Overall, our angle-resolved transport measurements in RHG establish an unprecedented scenario where rotational, time-reversal, and inversion symmetries are broken in tandem. These findings provide strong experimental support for the Coulomb-driven phenomenon of spontaneous momentum-space condensation  Huang et al. (2023); Jung et al. (2015); Dong et al. (2023); Parra-Martinez et al. (2025). According to prior theoretical discussions, momentum polarization could arise from the instability of an annular Fermi sea Huang et al. (2023); Jung et al. (2015), or from exchange-driven condensation between discrete Fermi pockets Dong et al. (2023); Parra-Martinez et al. (2025). These mechanisms provide a compelling framework for understanding the interplay between Fermi surface topology, isospin physics, and emergent ferroic orders in rhombohedral graphene systems.
acknowledgments
J.I.A.L. wishes to acknowledge helpful discussion with Chunli Huang and Rafael Fernandes. E.M. and J.I.A.L. acknowledge support from U.S. National Science Foundation under Award DMR-2143384. J.I.A.L. acknowledge partial support from the Air Force Office of Scientific Research. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. Part of this work was enabled by the use of pyscan (github.com/sandialabs/pyscan), scientific measurement software made available by the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy.